OpenStudy (anonymous):

Divide twelve and three-fifths divided by four and two-thirds. two and seven-tenths three and nine-tenths ten twenty-sevenths three and three-tenths

hartnn (hartnn):



hartnn (hartnn):

can you simplify ?

OpenStudy (anonymous):

is it, Divide, twelve and three-fifths divided by four and two-thirds. (with a comma after divide, like the first word just says that it is a division problem?)

hartnn (hartnn):

yes, correct

OpenStudy (anonymous):

okay, im here

OpenStudy (anonymous):

@carolinagb02, \[D\frac{a}{b}=\frac{(D \times b)+a}{b}\] \[\frac{a}{b} \div \frac{d}{c}= \frac{a}{b} \times \frac{c}{d}\]